OpenStudy (anonymous):

What is the 41st term of the arithmetic sequence where a1 = 17 and a15 = -39 ?

OpenStudy (anonymous):

@mathstudent55

OpenStudy (campbell_st):

well you need to find the common difference... which is the amount the sequence changes by the formula for a term in an arithmetic sequence is \[T_{n} = a + (n - 1)\times d\] n = number of terms, a = 1st term and d = common difference. to to find the common difference you know the 15th term, n = 15 and a = 17 substitute them to find d \[-39 = 17 +(15 -1)\times d\] or -39 = 17 + 14d solve for d. when you have it then use n = 41, a = 17 and your value of d in the formula to find the 41st term

OpenStudy (anonymous):

d is -4

OpenStudy (anonymous):

the Tn formula?

OpenStudy (campbell_st):

correct... now just go back to the formula to find the 41st term a = 17 and n = 41

OpenStudy (anonymous):